The effectiveness of a blood-pressure drug is being investigated. An experimenter finds that, on average, the reduction in systolic blood pressure is 68.8 for a sample of size 27 and standard deviation 7.4. Estimate how much the drug will lower a typical patient's systolic blood pressure (using a 99% confidence level). Assume the data is from a normally distributed population. Enter your answer as a tri-linear inequality accurate to three decimal places.

____ < μ < ____
Answer #1

sample mean = 68.8

n = 27

sd = 7.4

t value for 99% CI = TINV( 0.01, 26) = 2.779

E = t * \frac{s}{\sqrt{n}}

E = 2.779 * \frac{7.4}{\sqrt{27}}= 3.957

CI = mean +/- E = 68.8 +/- 3.957

CI = 64.843 , 72.757